find b if you happen to know the maximum y-value of the function y= -4x^2 + bx -1 is 21.
Answer by stanbon(75887)

find b if you happen to know the maximum y-value of the function
y= -4x^2 + bx -1 is 21
----------------
Comment:
Maximum occurs where x = -b/(2a) = -b/(-8) = b/8
-----
Max = f(b/8) = -4(b/8)^2 + b(b/8) -1 = 21
-----------
-4(b^2/64) + b^2/8 = 22
------
-b^2/16 +2(b^2/16) = 22
------
-b^2 + 2b^2 = 16*22
---
b^2 = 16*22
---------
b = 4*sqrt(22)
=====================
======================
Cheers,
Stan H.
============